# Cloud gap-filling with deep learning for improved grassland monitoring

### v1

@@ Abstract
Uninterrupted optical image time series are crucial for the timely monitoring of agricultural land changes. However, the continuity of such time series is often disrupted by clouds. In response to this challenge, we propose a deep learning method that integrates cloud-free optical (Sentinel-2) observations and weather-independent (Sentinel-1) Synthetic Aperture Radar (SAR) data, using a combined Convolutional Neural Network (CNN)-Recurrent Neural Network (RNN) architecture to generate continuous Normalized Difference Vegetation Index (NDVI) time series. We emphasize the significance of observation continuity by assessing the impact of the generated time series on the detection of grassland mowing events. We focus on Lithuania, a country characterized by extensive cloud coverage, and compare our approach with alternative interpolation techniques (i.e., linear, Akima, quadratic). Our method surpasses these techniques, with an average MAE of 0.024 and $\mathrm{R^{2}}$ of 0.92. It not only improves the accuracy of event detection tasks by employing a continuous time series, but also effectively filters out sudden shifts and noise originating from cloudy observations that cloud masks often fail to detect. A part of the code is accessible on Github.
@@ 1 Introduction
Grasslands cover around one third of the global agricultural land and they play a key role in food production, animal biodiversity and global carbon cycle (mara2012, ). They provide a wide range of ecosystem services, such as (i) provision of fodder for livestock, (ii) wildlife habitats, (iii) filtering or retention functions of waterways and greenhouse gas emissions, (iv) carbon storage, (v) pest control, (vi) crop pollination and (vii) protection against soil erosion (zhao2020grassland, ). The 2013 reform of the Common Agricultural Policy (CAP) has emphasized on the importance of grassland use intensity by requiring a minimum number of mowing or grazing events. National paying agencies have been tasked with the monitoring of compliance with this minimum activity condition. This information is important for the consolidation of agro-ecological measures and conservation schemes with regards to climate change (d'Adrimont2018, ; reinermann2023, ).
@@
The ever-increasing availability of remote sensing data has enabled the development of numerous competent agricultural monitoring systems. Most studies on grassland monitoring use remote sensing to address the problem of discriminating the different grassland categories (based on land cover, use intensity, etc.) and evaluate the various management strategies (ali2016, ). However, there are a few that focus on the quantification of grassland grazing intensity and detection of mowing events. In most cases, optical Earth observation data (MODIS, RapidEye, Landsat 8, Sentinel-2, etc.) are exploited (gomez2017, ; estel2018, ). A number of studies evaluate the mapping of the mowing or grazing frequency using statistical methods over multi-spectral image time series on regional or national scale (kolecka2018, ; Griffiths2019, ). Therefore, relevant vegetation indices, such as the Normalized Difference Vegetation Index (NDVI), the Leaf Area Index (LAI) or others, are used to detect drastic changes in the grasslands vegetation cover, between consecutive observations, which is related to the amount of biomass removed.
@@
Nevertheless, one of the main limitations of optical data is missing or invalid values caused by cloud coverage and unfavorable weather conditions. This problem is more evident in northern countries, where even the short revisit time of 5 days of Sentinel-2 can result in time series with long gaps (sudmanns2019, ). Consequently, frequent cloud cover in these areas obstructs the detection of mowing practices, rendering algorithms reliant on optical data less effective in discerning these events (kolecka2018, ). Hence, there is a need to artificially reconstruct these missing or corrupted values. This reconstruction can be performed by using only optical data, only SAR data or a collaborative exploitation of both (li2022, ). A dynamic combination of SAR with the available optical data can i) enrich the feature space and enhance the overall performance of downstream tasks and ii) fill in the missing values in optical products (garioud2021, ).
@@
The most straightforward method for addressing gaps in optical data is through temporal gap-filling techniques that rely only on optical data, such as the linear (or splines) interpolation (zeng2013, ). Despite the fact that these methods have been proved sufficient for several tasks (fauvel2019, ; sitok2021, ), in cases that we need to precisely track crops phenology development or identify any abrupt changes (e.g., detection of an event, natural disasters etc.), crucial points may be left out, and as a result, the performance would be sub-optimal, especially when the gaps are long (moreno2020, ).
@@
Similarly, the combination of various multi-sensor data can result in more dense satellite image time series (e.g., fusion of PlanetScope and Sentinel-2) which could, in turn, increase the number of cloud-free observations. However, there are still challenges that need to be taken into account, such as the different scales of spatial resolution (zhu2016, ) and the potential high financial cost of other sources (sadeh2021, ) that acts as a barrier to the scalability of the methodologies. Additionally, complementary inputs from ancillary optical sensors usually require numerous pre-processing steps to achieve the same data distribution (claverie2018, ) and the provision of sufficient number of additional observations is at stake since they are affected by cloud coverage as well.
@@
Contrary to the optical systems, SAR sensors can offer imagery at any time and under any weather circumstances, making them a very important source of information for agricultural monitoring in areas with frequent clouds (liu2019research, ). Moreover, backscatter coefficients can detect the geometric, structural and dielectric properties of plants (mandal2020dual, ), rendering SAR data suitable for tasks such as crop classification (ioannidou2022, ) and grassland monitoring (voormansik2016, ).
@@
In this study, we utilise Sentinel-1 and Sentinel-2 time series data to optimize the filling of missing NDVI values. Our approach involves a deep learning architecture that combines Convolutional and Recurrent Neural Networks (CNN-RNN). Specifically, we integrate Sentinel-1 backscatter and coherence time series with the available Sentinel-2 NDVI images to construct continuous NDVI time series of a fixed 6-day temporal resolution, focusing on grasslands across Lithuania. Then, we evaluate the impact of our generated continuous grassland time series in the downstream task of detecting grassland mowing. We choose the CNN-RNN architecture as it has been applied in the similar study of zhao2020 . The main contributions of this study are summarized as follows:
@@
    Evaluated the added-value of utilizing the continuous NDVI time series on detecting grassland mowing events.
@@ 2 Related work
The reconstruction of optical images has been widely studied, with most papers focusing on spectral, temporal, or spectral-temporal methods. Among these, spectral-temporal methods offer the best performance (shen2015missing, ). Recently, there has been an increase in studies that specifically target NDVI reconstruction. The methods used by these studies can be split into three main groups; i) temporal-based, ii) frequency-based, and iii) hybrid (li2021high, ). When it comes to vegetation monitoring, the most common are the temporal-based methods. These methods exploit the phenological continuity of NDVI time series, but they lack precision when long temporal gaps are present. Frequency-based methods have low computational complexity and are easy to implement, however, they are not able to handle sudden changes in vegetation growth, such as mowing events. On the other hand, hybrid methods, which utilize both temporal and spatial information, can achieve better results and are more capable of tackling the aforementioned issues. In the publication of julien2019optimizing , the authors evaluated and explored different parameterizations of five temporal reconstruction techniques. Even though the performance of some models was quite promising, the deviations in cloud-prone areas remained high.
@@
The methods mentioned above focus on filling missing NDVI values using only the available cloud-free time series. However, utilizing ancillary data sources, such as SAR data, provides valuable information that can be correlated with optical-derived vegetation indices in order to tackle the problem. In recent years, the extensive adoption of Artificial Intelligence (AI) has led to a rise in the utilization of Machine Learning (ML) approaches for gap-filling (wang2019, ; garioud2021, ). These approaches have shown promising performance across various land covers, including crops, grasslands, and forests. Unlike empirical models, which rely solely on initial hyperparameters and thresholds, these fully data-driven approaches require a large number of annotated training and validation samples to achieve high performance and ensure transferability and generalization.
@@
In literature, one can find many studies on the joint exploitation of Sentinel-1 and Sentinel-2 observations. Standard ML regression routines, such as random forest, support vector machines or gaussian processes have been applied to capture the correlations between the different sources and generate optical data, based exclusively on SAR observations (pipia2019, ; pinto2022, ). However, these approaches do not consider the temporal information of the input data and have proven to be very sensitive to the initial hyperparameter selection, depending on the type of vegetation cover and area characteristics, which challenges their generalization and robustness.
@@
The recent advancement in Deep Learning (DL), combined with increased computational capacities and the availability of open-access large datasets, paved the way for the successful adoption of fully automated data-driven approaches. Several works have been published that use CNN or Generative Adversarial Neural Networks (GANs) to exploit spatial characteristics and transform single date SAR images into equivalent optical ones (reyes2019, ; meraner2020, ). However, these approaches do not consider the dimension of time either, which is valuable information for crop monitoring, and usually have high computational complexity. On the other hand, RNNs are considered ideal for time series estimation because they are able to track temporal dependencies. Several architectures have been developed to exploit spatio-temporal data and enhance the prediction performance (rusvurm2018, ; ienco2019, ). Apart from RNNs, similar architectures can also effectively handle the cloud coverage issue, such as the one presented by turkoglu2021 .
In recent years, deep learning techniques that incorporate the temporal dimension have been effectively used for reconstructing optical vegetation indices using only SAR data. In zhao2020 , the authors propose the MCNN-seq framework which is a combination of a CNN and an LSTM layer. The CNN layer is used to reduce the impact of input noise and the LSTM layer, with an additional attention mechanism, addresses the problem of long time series to capture long-term dependencies in the data. However, despite this promising architecture, its generalization capability should be further assessed, since only a small number of crop classes and a limited geographical extent were evaluated.
In garioud2021 the authors propose a more sophisticated deep RNN design named SenRVM. This design involves the in-parallel processing of both descending and ascending orbits of Sentinel-1 SAR and coherence time series along with several auxiliary information, such as topographic, climatic and precipitation data. The approach’s capabilities are demonstrated through evaluation on three main vegetation types (grasslands, crops and forests) across two areas in France. The results suggest a satisfactory overall performance and emphasize its potential for generalization to large geographical areas. However, scaling up this methodology to the national level is difficult due to the challenge of acquiring the necessary auxiliary data and the time-consuming estimation of coherence coefficients. Nevertheless, for grasslands, the study found that ancillary information did not contribute significantly.
Recently, li2022 applied a Transformer Temporal-Spatial Model (TTSM) on a megapixel spatial scale (i.e., 1M pixels), which exhibited quite promising scores and outperformed other deep learning solutions, such as LSTM, CNN/LSTM and LSTM with attention. Here, the authors utilized a transfer learning approach to test their model on two neighboring areas in China. The one area was used for training, and the other was used for testing on the reconstruction of missing images for two vegetation types (crops and forests). Although the results are promising, it’s worth noting that the areas tested are geographically close to each other and share similar land cover characteristics.
Overall, the aforementioned solutions focus on the general problem of reconstructing NDVI using SAR data for multiple land cover types. However, to the best of our knowledge, no studies have explicitly addressed the impact of cloudy images on the detection of mowing events in grasslands.
@@
Grassland monitoring with remote sensing data has been widely studied. In order to detect grassland mowing, optical indices have been linked with interferometric SAR coherence of X and C bands (ali2017, ). However, confounding factors such as grazing events, farming activities, land factors and meteorological conditions can affect the accuracy of the results (tamm2016, ). Combining optical and radar data can enable the tracking of sudden events and mapping of grassland use intensity (reinermann2022detection, ). Several current methods use fixed thresholds after parameters calibration, leading to model instabilities, in terms of performance (garioud2021, ). The rule-based Sen4CAP algorithm (devroey2022, ) that uses both Sentinel-1 and Sentinel-2 time series was evaluated on six sites across Europe, achieving an overall accuracy of 69%. However, individual metrics on each area varied greatly (up to 30%).
@@
Several studies have used machine learning for mowing event detection. In particular, lobert2021 applied a deep learning model on sliding-windows of Sentinel-1/2 and Landsat 8 time series data for three consecutive years (2017-2019) in three regions of Germany. However, the study identified limitations, such as the interpolation of optical data to Sentinel-1 intervals, which may lead to the omission of events. Lange2022 developed a CNN-based algorithm to quantify land-use intensity over grasslands across Germany using Sentinel-2 NDVI time series. Finally, Holtgrave2023 explored machine learning potentials within a two-step transferable approach on initial optical time series gap-filling and a subsequent mowing event classification that combines optical, SAR, and weather data.
@@ 3.1 Study area
Lithuania (Fig. 1) is located in Northeastern Europe and covers an area of 65,300 km². Its temperate climate is influenced by both maritime and continental factors, and it is primarily defined as humid continental under the Köppen classification scheme. However, a small coastal zone adjacent to the Baltic Sea can be considered oceanic. It is mainly characterized by cold temperatures, frequent rainfalls, and snowy winters. As a result, four climatic regions have been identified: Coastal, Samogitian, Middle Lowlands, and Southeastern Highlands (lithuania_climatic_regions, ).
Grasslands are the most prevalent crop types and are spread equally across the country. They are divided into two main categories: permanent and temporary grasslands. Permanent are mostly perennial pastures and meadows, or other natural/semi-natural grassland parcels of the same land use for more than five years. On the other hand, temporary grasslands are the cases of newly updated grassland usage, of less than five years. In Lithuania, CAP requires farmers to perform at least one agricultural activity (mowing or grazing), usually until 1st of August of every cultivation period.
In this study, six distinct sites were selected based on their different agro-climatic and morphological characteristics as determined by Lithuania’s climatic region map and the distribution of grassland cultivation.
Region 1 is located in the county of Alytus (southern Lithuania) at the center of the Dzūkija region that is mostly covered with pine forests. Despite being watered by the Nemunas river, this area is unsuitable for agricultural activities due to the poor quality of the soil, which is mostly sandy. Regions 3 and 5 are located around the counties of Utena and Panevėžys, respectively, in the northeast part of the country and inside the region of Aukštaitija, which contains more urban elements than the rest. In the Samogitia region, an area in the northwestern Lithuania with many lakes and intense agricultural activity, we find Regions 4 and 6, near the counties of Telšiai and Šiauliai, respectively. Finally, Region 2 is in the county of Klaipėda, located in the western part of the country. Its climate is highly influenced by the Baltic Sea, with very frequent storms, as well as snowfalls and gales during winter. Table 1 contains the number of grassland parcels and their average size for each study site. The average size of grassland parcels in entire country is over 1 ha (see the respective Figure S1 in A), which is substantial bigger in comparison to other European countries. However, there are a few cases with size less than 0.1 ha, which results to mixed information within the Sentinel-2 pixels and have been eliminated from our study.
Table 1: Number of grasslands parcels and their respective average size in hectares (ha) for the six study regions.

### v2

@@ Abstract
Uninterrupted optical image time series are crucial for the timely monitoring of agricultural land changes, particularly in grasslands. However, the continuity of such time series is often disrupted by clouds. In response to this challenge, we propose an innovative deep learning method that integrates cloud-free optical (Sentinel-2) observations and weather-independent (Sentinel-1) Synthetic Aperture Radar (SAR) data. Our approach employs a hybrid architecture combining Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) to generate continuous Normalized Difference Vegetation Index (NDVI) time series, highlighting the role of NDVI in the synergy between SAR and optical data. We demonstrate the significance of observation continuity by assessing the impact of the generated NDVI time series on the downstream task of grassland mowing event detection. We conducted our study in Lithuania, a country characterized by extensive cloud coverage, and compared our approach with alternative interpolation techniques (i.e., linear, Akima, quadratic). Our method outperformed these techniques, achieving an average Mean Absolute Error (MAE) of 0.024 and a coefficient of determination ($\mathrm{R^{2}}$) of 0.92. Additionally, our analysis revealed improvement in the performance of the mowing event detection, with F1-score up to 84% using two widely applied mowing detection methodologies. Our method also effectively mitigated sudden shifts and noise originating from cloudy observations, which are often missed by conventional cloud masks and adversely affect mowing detection precision.
@@ 1 Introduction
Grasslands cover around one-third of the global agricultural land and they play a key role in food production, animal biodiversity, and global carbon cycle [1]. They provide a wide range of ecosystem services, such as (i) provision of fodder for livestock, (ii) wildlife habitats, (iii) filtering or retention functions of waterways and greenhouse gas emissions, (iv) carbon storage, (v) pest control, (vi) crop pollination and (vii) protection against soil erosion [2]. The 2013 reform of the Common Agricultural Policy (CAP) has emphasized the importance of grassland use intensity by requiring a minimum number of mowing or grazing events. Therefore, national paying agencies have been tasked with monitoring compliance with this minimum activity condition. This information is important for the consolidation of agro-ecological measures and conservation schemes with regards to climate change [3, 4].
@@
The ever-increasing availability of remote sensing data has enabled the development of numerous competent agricultural monitoring systems. Most studies on grassland monitoring use remote sensing to address the problem of discriminating the different grassland categories (based on land cover, use intensity, etc.) and evaluate the various management strategies [5, 6]. However, few focus on the quantification of grassland grazing intensity and the detection of mowing events. In most cases, optical Earth observation data (MODIS, RapidEye, Landsat 8, Sentinel-2, etc.) are exploited [7, 8]. Several studies evaluate the mapping of the mowing or grazing frequency using statistical methods over multi-spectral image time series on a regional or national scale [9, 10, 11]. Therefore, relevant vegetation indices, such as the Normalized Difference Vegetation Index (NDVI), the Leaf Area Index (LAI), or others, are used to detect drastic changes in the grassland vegetation cover, between consecutive observations, which is related to the amount of biomass removed.
@@
Nevertheless, one of the main limitations of optical data is missing or invalid values caused by cloud coverage and unfavorable weather conditions. This problem is more evident in northern countries, where even the short revisit time of 5 days of Sentinel-2 can result in time series with long gaps [12]. Consequently, frequent cloud cover in these areas obstructs the detection of mowing practices, rendering algorithms reliant on optical data less effective in discerning these events [9]. Hence, there is a need to artificially reconstruct these missing or corrupted values. This reconstruction can be performed by using only optical data, only Synthetic Aperture Radar (SAR) data, or collaborative exploitation of both [13].
@@
The most straightforward method for addressing gaps in optical data is to use temporal gap-filling techniques that rely only on optical data, such as the linear (or splines) interpolation [14]. Even though these methods have proved sufficient for several tasks [15, 16], in cases where we need to precisely track crop phenology development [17, 18] or identify any abrupt changes (e.g., cultivation practices, natural disasters etc.) [19], crucial points may be left out, and as a result, the performance would be sub-optimal, especially when the gaps are long [20].
@@
Equally, the combination of various multi-sensor data can result in more dense satellite image time series (e.g., fusion of PlanetScope and Sentinel-2) which could, in turn, increase the number of cloud-free observations. However, there are still challenges that need to be taken into account, such as the different scales of spatial resolution [21] and the potentially high financial cost of other sources [22] that acts as a barrier to the scalability of the methodologies. Additionally, complementary inputs from ancillary optical sensors usually require numerous pre-processing steps to achieve the same data distribution [23], and the provision of a sufficient number of additional observations is at stake since they are affected by cloud coverage as well.
@@
Contrary to the optical systems, SAR sensors can offer imagery under any weather circumstances, making them a very important source of information for agricultural monitoring in areas with frequent clouds [24]. Moreover, backscatter coefficients can detect the geometric, structural, and dielectric properties of plants [25], rendering SAR data suitable for tasks such as crop classification [26] and grassland monitoring [27]. Overall, a dynamic combination of SAR with the available optical data can i) enrich the feature space and enhance the overall performance of downstream tasks and ii) fill in the missing values in optical products [28].
@@
In this study, we utilize Sentinel-1 and Sentinel-2 time series data to optimize the filling of missing NDVI values. Our approach involves a deep learning architecture that combines Convolutional and Recurrent Neural Networks (CNN-RNN). Specifically, we integrate Sentinel-1 backscatter and coherence time series with the available Sentinel-2 NDVI images to construct continuous NDVI time series of a fixed 6-day temporal resolution, focusing on grasslands across Lithuania. Then, we evaluate the impact of our generated continuous grassland time series in the downstream task of detecting grassland mowing activity. We choose the CNN-RNN architecture as it has been applied in a similar study of [29]. The main contributions of this study are summarized as follows:
@@
    Evaluated the added value of utilizing the continuous NDVI time series on detecting grassland mowing events.
@@ 2 Related work
The reconstruction of optical images has been widely studied, with most papers focusing on spectral, temporal, or spectral-temporal methods. Among these, spectral-temporal methods offer the best performance [30]. Recently, there has been an increase in studies that specifically target NDVI reconstruction. The methods used by these studies can be divided into three main groups: i) temporal-based, ii) frequency-based, and iii) hybrid [31]. When it comes to vegetation monitoring, the most common are the temporal-based methods [32]. These methods exploit the phenological continuity of NDVI time series, but they lack precision when long temporal gaps are present. Frequency-based methods [33] have low computational complexity and are easy to implement, however, they are not able to handle sudden changes in vegetation growth, such as mowing events. On the other hand, hybrid methods [34, 35], which utilize both temporal and spatial information, can achieve better results and are more capable of tackling the aforementioned issues. In the publication of [36], the authors evaluated and explored different parameterizations of five temporal reconstruction techniques. Even though the performance of some models was quite promising, the deviations in cloud-prone areas remained high.
@@
The abovementioned methods focus on filling in missing NDVI values using only the available cloud-free time series. However, utilizing ancillary data sources, such as SAR data, provides valuable information that can be correlated with optical-derived vegetation indices. In recent years, the extensive adoption of Artificial Intelligence (AI) has led to a rise in the utilization of Machine Learning (ML) approaches for gap-filling [37, 28]. These approaches have shown promising performance across various land covers, including crops, grasslands, and forests. Unlike empirical models, which rely solely on initial hyperparameters and thresholds, these fully data-driven approaches require a large number of annotated training and validation samples to achieve high performance and ensure transferability and generalization.
@@
In literature, one can find many studies on the joint exploitation of Sentinel-1 and Sentinel-2 observations [38, 39, 40]. For instance, standard ML regression routines have been applied to capture the correlations between the different sources and generate optical data, based exclusively on SAR observations [41]. Additionally, [42] demonstrated the effectiveness of random forest in combining Sentinel-1 and Landsat-8 satellite data to predict vegetation indices accurately. Similarly, [43] explored Gaussian processes to model the relationship between SAR and optical data, emphasizing the method’s robustness in filling cloud-induced data gaps. However, these approaches do not consider the temporal dimension of the input data and have proven to be very sensitive to the initial hyperparameter selection, depending on the type of vegetation cover and area characteristics, which challenges their generalization and overall effectiveness.
@@
The recent advancement in Deep Learning (DL), combined with increased computational capacities and the availability of open-access large datasets, paved the way for the successful adoption of fully automated data-driven approaches. Several works have been published that use CNN or Generative Adversarial Neural Networks (GANs) to exploit spatial characteristics and transform single-date SAR images into equivalent optical ones [44, 45, 46]. However, these approaches do not consider the dimension of time either, which is valuable information for crop monitoring, and usually have high computational complexity. On the other hand, RNNs are considered ideal for time series estimation because they are able to track temporal dependencies. Several architectures have been developed to exploit spatio-temporal data and enhance the prediction performance [47, 48]. Interestingly, [49] found that certain RNN cells evidenced sensitivity to cloudy pixels, potentially acting as an internal filtering mechanism.
Over the past several years, deep learning techniques that incorporate the temporal dimension have been effectively used for reconstructing optical vegetation indices using only SAR data. In [29], the authors proposed the MCNN-seq framework, combining CNN and Long Short-Term Memory (LSTM) layers with an attention mechanism to capture long-term dependencies in the time series. However, its generalization capability needs further assessment due to the small number of crop classes and the limited geographical extent (i.e., Southern California). In [28] the authors introduced SenRVM, a deep RNN design that processes Sentinel-1 SAR data alongside auxiliary information such as topographic and climatic details. This approach showed promising performance across various vegetation types in France, but scaling up to a national level poses challenges due to the extensive data requirements and computational complexity. No less important, [13] applied a Transformer Temporal-Spatial Model (TTSM) on a larger spatial scale (i.e., 1M pixels), outperforming other deep learning architectures. They used transfer learning to test their model on neighboring areas in China, achieving promising results for crops and forests. However, the geographical proximity of the test regions suggests limited generalizability. Lately, [50] proposed a novel methodology that combines sparse optical NDVI time series with denser but less accurate SAR-derived NDVI time series using a Gated Recurrent Unit (GRU), applied on 1206 regions worldwide between 2019 and 2021. Overall, the aforementioned solutions focus on the general problem of reconstructing NDVI from SAR data across various land cover types. Yet, to the best of our knowledge, no studies have explicitly addressed the impact of cloudy images on the detection of mowing events in grasslands.
@@
Grassland monitoring with remote sensing data has been widely studied. In order to detect grassland mowing, optical indices have been linked with interferometric SAR coherence of X and C bands [51, 52]. However, confounding factors such as grazing events or other farming activities, land factors and meteorological conditions can affect the accuracy of the results [53]. To this purpose, combining optical and radar data can enable the tracking of sudden events and mapping of grasslands use intensity [54].
Current methods using fixed thresholds can lead to performance instabilities [28]. The rule-based Sen4CAP algorithm [55] includes a Sentinel-2 component in its processing chain for mowing detection, which searches for sharp drops between consecutive NDVI cloud-free observations, surpassing a predefined threshold. The algorithm achieved 69% accuracy across six European sites, though individual site metrics varied significantly (up to 30%). In 2019, it demonstrated an F1-score of 79% over 150 parcels in the Netherlands. However, when evaluated on 118 cases in Lithuania, the F1-score dropped to 60%, primarily due to a significantly reduced precision of 49%. Another widely used approach is the algorithm proposed by [56]. This method considers the residuals of the assessed vegetation index values from an interpolated time series formulated based on the seasonal peak timestamps that represent the ideal growing season curve. Indicatively, the algorithm achieved an overall F1-score of 67% on 180 grassland parcels in Germany for 2020.
@@
Similarly, several studies have used machine learning for mowing event detection [57, 58]. In particular, [59] applied a deep learning model on Sentinel-1/2 and Landsat 8 time series data in Germany, identifying limitations such as potential event omission due to optical data interpolation to Sentinel-1 intervals. [60] developed a CNN-based algorithm to quantify land-use intensity over grasslands using Sentinel-2 NDVI time series. Finally, [61] explored a two-step transferable approach that initially uses optical, SAR, and weather data to fill gaps in optical time series, followed by a binary classification to distinguish between mown and unmown grassland cases.
@@ 3.1 Study area
Lithuania is located in Northeastern Europe and covers an area of 65,300 km². Its temperate climate is influenced by both maritime and continental factors, and it is primarily defined as humid continental under the Köppen classification scheme. However, a small coastal zone adjacent to the Baltic Sea can be considered oceanic. It is mainly characterized by cold temperatures, frequent rainfalls, and snowy winters. As a result, four climatic regions have been identified: Coastal, Samogitian, Middle Lowlands, and Southeastern Highlands [62].
Grasslands are the most prevalent crop types and are spread equally across the country. They are divided into two main categories: permanent and temporary grasslands. Permanent are mostly perennial pastures and meadows, or other natural/semi-natural grassland parcels of the same land use for more than five years. On the other hand, temporary grasslands are the cases of newly updated grassland usage, of less than five years. In Lithuania, CAP requires farmers to perform at least one agricultural activity (mowing or grazing), usually until the 1st of August of every cultivation period.
In this study, six distinct regions were selected based on their different agro-climatic and morphological characteristics as determined by Lithuania’s climatic region map and the distribution of grassland cultivation (see Fig. 1).
Figure 1: Study regions in Lithuania.
Region 1 is located in the county of Alytus (southern Lithuania) at the center of the Dzūkija region which is mostly covered with pine forests. Despite being watered by the Nemunas River, this area is unsuitable for agricultural activities due to the poor quality of the soil, which is mostly sandy. Regions 3 and 5 are located around the counties of Utena and Panevėžys, respectively, in the northeast part of the country and inside the region of Aukštaitija, which contains more urban elements than the rest. In the Samogitia region, an area in northwestern Lithuania with many lakes and intense agricultural activity, we find Regions 4 and 6, near the counties of Telšiai and Šiauliai, respectively. Finally, Region 2 is in the county of Klaipėda, located in the western part of the country. Its climate is highly influenced by the Baltic Sea, with very frequent storms, as well as snowfalls and gales during winter. Table 1 contains the number of grassland parcels and their average size for each study region. The average size of grassland parcels in the entire country is over 1 ha (see the respective Fig. S1 in A), which is substantially bigger in comparison to other European countries. However, there are a few cases with size less than 0.1 ha, which results in mixed information within the Sentinel-2 pixels and have been eliminated from our study.
Table 1: Number of grassland parcels and their respective average size in hectares (ha) for the six study regions.